# Zero-Shot Statistical Tests for LLM-Generated Text Detection using Finite Sample Concentration Inequalities

### v3

@@ Appendix B Related work
With the advent of LLMs, human author attribution extends to (1) human-written vs. machine-generated detection, (2) machine-generated vs. another machine-generated attribution, and (3) attribution in text generated through LLM and human collaboration. These shifts necessitate transitioning from human-centric stylometry to detection frameworks tailored for LLMs. For a survey that connects the older literature on authorship attribution to modern problems in the era of LLMs, we refer readers to Huang et al. (2025).
Statistical tests on LLM-generated text. Recent efforts have provided statistical tests for detecting LLM-generated text and determining whether cloud implementation works the same as the reference model. For example, Li et al. leverages the watermark key sequence and next token probabilities (NPTs) to introduce statistical tests for detecting LLM-generated text when the detection approach is watermarking, and their method applies to short and long text. Gao et al. (2024) introduces a framework for detecting whether an API provider has modified a language model’s output distribution without informing users. The paper formalizes the problem as a two-sample statistical testing task, where users collect samples from an API and compare them to a reference model’s outputs. They propose a hypothesis testing method using Maximum Mean Discrepancy (MMD) with a string kernel, first proposed in Gretton et al. (2012), to detect whether the API serves the expected model or a modified version.

### v4

@@ Appendix C Literature review
With the advent of LLMs, human author attribution extends to (1) human-written vs. machine-generated detection, (2) machine-generated vs. another machine-generated attribution, and (3) attribution in text generated through LLM and human collaboration. These shifts necessitate transitioning from human-centric stylometry to detection frameworks tailored for LLMs. For a survey that connects the older literature on authorship attribution to modern problems in the era of LLMs, we refer readers to Huang et al. [26].
Statistical tests on LLM-generated text.
Recent research has introduced statistical testing for detecting LLM-generated text and verifying whether the output behavior of deployed (e.g., cloud-hosted) models remains consistent with a known reference model. Li et al. [39], focuses on detecting watermarked text. Their method uses the next-token probabilities (NPTs) to construct likelihood-based hypothesis tests. Gao et al. [15] determines whether a commercial or third-party API provider has modified the underlying language mode, which is a practical concern when users rely on cloud-accessed LLMs whose internals may be updated without notice. They formalize this as a two-sample hypothesis testing problem, where users compare outputs from the deployed API against samples from a previously archived reference model. To quantify distributional shifts, they propose using Maximum Mean Discrepancy (MMD) with a string kernel, a nonparametric method first introduced in Gretton et al. [19], which enables detection of subtle changes in the model’s token distribution. Our framework is zero-shot, and as a result, it requires neither labeled LLM outputs nor access to the generation model (as elaborated in Section 4.2), and we also theoretically show that the type I and type II errors of our test decrease exponentially in the text length.
Asymptotic features of LLM-generated text.
Mudireddy et al. [43] shows that the log-perplexity of any large text generated by a language model must asymptotically converge to the average entropy of its token distributions when the evaluator and the generator model are the same. Their result relies on sufficiently large text samples to reach asymptotic behavior. This version of their manuscript, with correct proofs, was made online at the same time/after the initial release of our manuscript. Our work is a generalization to the case with an arbitrary length text and the case where generator and evaluator models are different by showing the convergence of log-perplexity to cross-entropy between the generator and the evaluator models, except with an exponentially small probability in the string length.